SOLUTION SET OF (3/5)x=(x -x2-9) is?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear vardaan

(3/5)x=(x -x2-9)

L.HS = (3/5)x  always positive

RHS = (x -x2-9) 

    coefficient of x2 is negative  and discrimnent is also negative so this expresion always has negative value for all x

 

so no solution exists.

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed solution very quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin